Circular No. 8119

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002km–2002ky M. Doi, University of Tokyo, on behalf of the Supernova Cosmology

Project (cf. IAUC 7971, plus V. Fadeyev, B. Lee, V. Stanishev, and R. Vogel) and the Subaru high-redshift supernova search group (N. Yasuda, N. Kashikawa, K. Motohara, T. Morokuma, K. Sekiguchi, G. Kosugi, H. Furusawa, Y. Komiyama, T. Takata, M. Ouchi, Y. Ohyama, and Subaru Observatory SXDS Project members), reports the discovery of 13 spectroscopically confirmed supernovae found with the Subaru telescope (+ Suprime-Cam) in the SDSS i′ band. None of the supernovae were present in reference images (limiting mag ∼ 26.6; S/N = 5 for 2′′ aperture) taken on 2002 Sept. 30 and Oct. 1. The supernovae spectra were obtained at the Gemini North telescope (+ GMOS) on Nov. 6, 8, and 9; at the Keck II telescope (+ ESI) on Nov. 6, 7, 9, and 10; at the Very Large Telescope Yepun UT4 (+ FORS2) on Nov. 7–11; and at the Subaru telescope (+ FOCAS) on Nov. 12. Redshifts were obtained using either the host galaxy spectrum (denoted in the table with an asterisk) or with template spectrum fitting of a supernova. SN 2002kr is in an elliptical host galaxy; it — together with SNe 2002km, 2002kn, and 2002ko — have spectra consistent with that of a type-Ia supernova. The spectra of the other objects show broad (several tens of nm wide) lines that are indicative of supernovae and allow the ruling out of active galactic nuclei, but the low signal-to-noise and limited wavelengths prevent definitive typing.

SN  2002 UT  α_{2000}  δ_{2000}  i′ z  Offset
2002km Nov. 3.3 2^{h}16^{m}23.^{s}93 − 4^{o}49′29′′.4 24.5 0.606* 0′′.2 W
2002kn Nov. 3.3 2 16 45.71 − 5 09 51.2 25.0 1.03 (no host)
2002ko Nov. 3.3 2 17 00.05 − 4 58 19.6 23.4 0.57 0′′.5 W
2002kp Nov. 3.3 2 17 08.63 − 5 02 06.4 23.8 0.928* 1′′.4 E
2002kq Nov. 3.3 2 17 12.24 − 4 55 08.7 24.4 0.823* 0′′.3 NW
2002kr Nov. 3.3 2 17 34.51 − 4 53 46.6 24.5 1.063* —
2002ks Nov. 3.3 2 17 34.53 − 5 00 15.4 25.2 1.181* 1′′.3 SE
2002kt Nov. 3.3 2 17 38.08 − 5 08 46.8 24.5 0.505* 0′′.3 NW
2002ku Nov. 3.3 2 17 42.54 − 5 06 34.0 24.8 0.92* 0′′.5 NE
2002kv Nov. 3.3 2 17 43.30 − 4 30 56.7 24.6 0.926* 0′′.4 E
2002kw Nov. 3.3 2 18 35.15 − 4 26 38.9 25.1 0.59 0′′.6 NW
2002kx Nov. 3.3 2 18 40.73 − 5 03 44.3 25.3 0.623* 1′′.1 NW
2002ky Nov. 3.3 2 18 53.20 − 4 32 59.2 23.7 0.66 0′′.6 N

2003 April 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green